# Huairou 101408,Beijing,China

### v1

@@ 1 Introduction and summary
We solve the gravitation 7-vertex model by mapping it to the large-$N$ limit of an $U(N)$-invariant matrix model. The matrix model in question is one-parameter generalisation of the 6-vertex matrix model studied in ZinnJustin:1999wt ; Kostov:1999qx ; ELVEYPRICE2023105739 . With the volume of the $U(N)$ symmetry group factored out, the partition function of the matrix model is an $N$-fold integral
@@
where the dependence on the temperature coupling is encoded in the cubic potential $V(x)$ and the parameter $q=-w^{2}$. In the large $N$ limit, the integral depends on two adjustable parameters, the temperature $T$ and the lattice cosmological constant $\kappa=\hbar N$. Interesting critical behaviour occurs only if $|q|=1$. For q=1q=-1, this is the partition function of the gravitational $O(2)$ model. We will parametrise the vertices in fig. 1 in a way compatible with the conventions of Kostov:1999qx by setting in (1)
@@
The free energy ${\cal F}=\hbar\log{\mathcal{Z}}$ takes simpler form in the grand canonical ensemble where the bulk cosmological constant $\mu$ is conjugated to the size $N$ of the matrix and a sum over $N$ is performed. We find that the boundary entropy $M$ and the susceptibility $u(\mu,t)=\partial_{\mu}^{2}{\cal F}$ are related by
@@
The plan of the paper is as follows. We start by reviewing in sect. 2 the phase structure of the sine-Gordon model and sine-Liouville gravity. In sect. 3 we define partition function of the vertex model and derive its representations in terms of a gas of oriented loops on planar graphs and as an interaction-round-a-triangle height model. Then we formulate the dual large-$N$ matrix model and formulate the Dyson-Schwinger equations as Virasoro constraint for a collective bosonic field.
@@ 2.2 Sine-Liouville gravity
$\begin{aligned} {\mathcal{A}}_{{}_{{SL}}}[\phi,\varphi]&={1\over 4\pi}\int\limits_{\mathcal{M}}d^{2}z\sqrt{\hat{g}}\left[(\hat{\nabla}\phi)^{2}+(\hat{\nabla}\varphi)^{2}Q\phi\hat{R}\right]+{\rm ghosts}\\ &+\mu\,\int d^{2}z\sqrt{\hat{g}}\,e^{2b\phi}+\sum_{\pm}{\mu_{\pm}}\int d^{2}z\sqrt{\hat{g}}\ e^{\alpha\phi}e^{\pm ia\varphi},\qquad a={1\over R}.\end{aligned}$ (21)
@@ 3.2 The 7-vertex matrix model
The gaussian integral over the complex matrix can be done explicitly, leaving an $\mathbf{SU}(N)$-invariant integral over the hermitian matrix $\mathbf{X}$ which reduces to an $N$-fold integral with respect to the real eigenvalues $x_{1},...,x_{N}$,
@@
The integrand in (3) is an $N\times N$ determinant and can be formulated as the partition function of a system of $N$ fermions. In view of comparing with the MQM approach to sine-Liouville gravity Alexandrov:2002fh , it is more convenient to work in the grand ensemble defined as
@@
with the Fredholm kernel defined with flat integration measure $dx$ on the real axis. The Fredholm determinant describes the grand ensemble of free fermions with fixed Fermi energy $\mu$.
@@
where the expectation value of the trace in the grand ensemble (32) is defined as the sum, with a weight $e^{-\mu N/\hbar}$, of the unnormalised expectation values in the sectors with $N$ eigenvalues. The reslvent of the random matrix
@@ 4.6 Computation of the boundary entropy
In the parametrisation strip, the now infinite $A$-cycle connects the points $s=$-$\infty$ with $s=+\infty$, and the $B$-cycle connects the two edges of the $s$-parametrisation strip. The second relation (55) states, in terms of $\varphi(s)\equiv\varphi[x(s)]$, that for any $s$,
@@ 4.7 Partition function on the sphere and susceptibility
Since the $A$-cycle becomes infinite in the continuum limit, we regularise the integral (77) by introducing an UV cutoff $\Lambda$. Retaining only the leading term in the large $x$ expansion, we have
@@ 6 Discussion
In this paper we proposed a microscopic realisation of the sine-Liouville gravity based on a “dilute” vertex model on dynamical lattice which represents a one-parameter generalisation of the six-vertex model considered earlier. We extracted the phase diagram of the theory from the classical spectral curve of the dual large-$N$ matrix model and gave a QFT interpretation of the critical points, helped by the phase diagram of the sine-Gordon model on a flat surface. In particular, we identified the gravitational equivalent of the massless flow in the sine-Gordon model with imaginary coupling constant.

### v2

@@ 1 Introduction and summary
We solve the gravitation 7-vertex model by mapping it to the large-NN limit of an $U(N)$-invariant matrix model. The matrix model in question is one-parameter generalisation of the 6-vertex matrix model studied in ZinnJustin:1999wt ; Kostov:1999qx ; ELVEYPRICE2023105739 . With the volume of the $U(N)$ symmetry group factored out, the partition function of the matrix model is an NN-fold integral
@@
where the dependence on the temperature coupling is encoded in the cubic potential $V(x)$ and the parameter $q=-w^{2}$. In the large NN limit, the integral depends on two adjustable parameters, the temperature $T$ and the lattice cosmological constant $\kappa=\hbar N$. Interesting critical behaviour occurs only if $|q|=1$. For q=1q=-1, this is the partition function of the gravitational $O(2)$ model. We will parametrise the vertices in fig. 1 in a way compatible with the conventions of Kostov:1999qx by setting in (1)
@@
The free energy ${\cal F}=\hbar\log{\mathcal{Z}}$ takes simpler form in the grand canonical ensemble where the bulk cosmological constant $\mu$ is conjugated to the size NN of the matrix and a sum over NN is performed. We find that the boundary entropy $M$ and the susceptibility $u(\mu,t)=\partial_{\mu}^{2}{\cal F}$ are related by
@@
The plan of the paper is as follows. We start by reviewing in sect. 2 the phase structure of the sine-Gordon model and sine-Liouville gravity. In sect. 3 we define partition function of the vertex model and derive its representations in terms of a gas of oriented loops on planar graphs and as an interaction-round-a-triangle height model. Then we formulate the dual large-NN matrix model and formulate the Dyson-Schwinger equations as Virasoro constraint for a collective bosonic field.
@@ 2.2 Sine-Liouville gravity
$\begin{aligned} {\mathcal{A}}_{{}_{{SL}}}[\phi,\varphi]&={1\over 4\pi}\int\limits_{\mathcal{M}}d^{2}z\sqrt{\hat{g}}\left[(\hat{\nabla}\phi)^{2}+(\hat{\nabla}\varphi)^{2}+Q\phi\hat{R}\right]+{\rm ghosts}\\ &+\mu\,\int d^{2}z\sqrt{\hat{g}}\,e^{2b\phi}+\sum_{\pm}{\mu_{\pm}}\int d^{2}z\sqrt{\hat{g}}\ e^{\alpha\phi}e^{\pm ia\varphi},\qquad a={1\over R}.\end{aligned}$ (21)
@@ 3.2 The 7-vertex matrix model
The gaussian integral over the complex matrix can be done explicitly, leaving an $\mathbf{SU}(N)$-invariant integral over the hermitian matrix $\mathbf{X}$ which reduces to an NN-fold integral with respect to the real eigenvalues $x_{1},...,x_{N}$,
@@
The integrand in (3) is an $N\times N$ determinant and can be formulated as the partition function of a system of NN fermions. In view of comparing with the MQM approach to sine-Liouville gravity Alexandrov:2002fh , it is more convenient to work in the grand ensemble defined as
@@
with the Fredholm kernel defined with flat integration measure dxdx on the real axis. The Fredholm determinant describes the grand ensemble of free fermions with fixed Fermi energy $\mu$.
@@
where the expectation value of the trace in the grand ensemble (32) is defined as the sum, with a weight $e^{-\mu N/\hbar}$, of the unnormalised expectation values in the sectors with NN eigenvalues. The reslvent of the random matrix
@@ 4.6 Computation of the boundary entropy
In the parametrisation strip, the now infinite AA-cycle connects the points $s=$-$\infty$ with $s=+\infty$, and the $B$-cycle connects the two edges of the $s$-parametrisation strip. The second relation (55) states, in terms of $\varphi(s)\equiv\varphi[x(s)]$, that for any $s$,
@@ 4.7 Partition function on the sphere and susceptibility
Since the AA-cycle becomes infinite in the continuum limit, we regularise the integral (77) by introducing an UV cutoff $\Lambda$. Retaining only the leading term in the large $x$ expansion, we have
@@ 6 Discussion
In this paper we proposed a microscopic realisation of the sine-Liouville gravity based on a “dilute” vertex model on dynamical lattice which represents a one-parameter generalisation of the six-vertex model considered earlier. We extracted the phase diagram of the theory from the classical spectral curve of the dual large-NN matrix model and gave a QFT interpretation of the critical points, helped by the phase diagram of the sine-Gordon model on a flat surface. In particular, we identified the gravitational equivalent of the massless flow in the sine-Gordon model with imaginary coupling constant.